Project L costs $38,531.91, its expected cash inflows are $9,000 per year for 8 years, and its WACC is 12%. What is the project's IRR? Round your answer to two decimal places.
Answer #1

Investment is 38531.91

Expected cash flows are 9000 for 8 years

Wacc is 12%

IRR is discount rate at which npv will become 0

Npv is present value of cashflows less initial investment

Npv at 12% discount rate

=9000(PVIFA 12% 8years) - 38531.91

= 9000(4.9676) -38531.91

= 6176.49

Npv at 13% discount rate

= 9000(PVIFA 13% 8 years) -38531.91

= 9000(4.7988) -38531.91

= 4676.29

IRR can be calculated by using interpolation method

Formula is

Lower rate +[ NPV at lower rate / (NPV at lower rate - NPV at higher rate)] * ( Higher rate - Lowe rate)

= 12% +(6176.49)/(6176.49-4676.29) = 16.11%